i really need to get my grade up in order to pass 9th grade !
write an equation in point-slope form of the line that passes through the given point and has the given slope.
(2,5), m=3

Found 2 solutions by lefty4ever26, amoresroy:
Answer by lefty4ever26(59)

y-y1=m(x-x1) you plug in (2,5) into y1 and x1 and your slope is m=3
SO,
y-5=3(x-2)
y-5=3x-6
y=3x-1 and this is the equation of the line.

Answer by amoresroy(361)

write an equation in point-slope form of the line that passes through the given point and has the given slope.
(2,5), m=3
point-slope form is: Y - Y1 = m(X - X1)
Y - 5 = 3(X - 2)